40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jun 3 16:46:31 2019, gautam, Update, IOO, IMC not locking PCdrive_RMS.png
    Reply  Tue Jun 4 10:56:31 2019, gautam, Update, IOO, IMC diagnostics IMCdiag.pdf
       Reply  Thu Jun 6 22:11:53 2019, Koji, Update, IOO, IMC diagnostics 
          Reply  Thu Jun 13 18:01:18 2019, aaron, Update, IOO, IMC diagnostics 
             Reply  Fri Jun 14 13:10:00 2019, aaron, Update, IOO, IMC diagnostics 
                Reply  Mon Jun 17 12:37:16 2019, aaron, Update, IOO, IMC diagnostics 
                   Reply  Tue Jun 18 20:35:07 2019, aaron, Update, IOO, IMC diagnostics 
                      Reply  Wed Jun 19 19:12:51 2019, aaron, Update, IOO, IMC diagnostics PMC_Error_Spectrum.pdfPDH_signal.pdfPDH_signal_full.pdf
                         Reply  Mon Jun 24 15:49:05 2019, aaron, Update, IOO, IMC diagnostics sweep_config_updated.png
                            Reply  Tue Jul 2 19:36:18 2019, aaron, Update, IOO, IMC diagnostics PDH_error.pdfPMC_Error_Spectrum.pdf
          Reply  Fri Jun 21 19:36:26 2019, Koji, Update, IOO, IMC diagnostics IMC_error.pdf
             Reply  Sun Jun 23 08:09:53 2019, gautam, Update, IOO, NPRO diagnostics consolidatedOscopeScreenCaps.pdfRINcomp.pdfPLL_OLTF.pdfPLLnoise.pdf
                Reply  Sun Jun 23 09:36:32 2019, gautam, Update, IOO, IMC is locking normally again PMCdiag.pdf
                   Reply  Sun Jun 23 14:43:14 2019, Koji, Update, IOO, IMC is locking normally again 
                   Reply  Mon Jun 24 08:12:10 2019, gautam, Update, IOO, IMC is locking normally again IMCdutycycle.png
                      Reply  Mon Jun 24 11:48:35 2019, gautam, Update, IOO, IMC in-loop error spectra and OLTF IMCfreqNoise.pdfIMC_OLTF.pdfIMC_CMboard.jpg
Message ID: 14687     Entry time: Sun Jun 23 08:09:53 2019     In reply to: 14686     Reply to this: 14688  
Author: gautam 
Type: Update 
Category: IOO 
Subject: NPRO diagnostics 

Summary:

Over the last few days, I've been doing some (complementary) measurements to what Aaron and Koji have been looking at. The motivation was to identify if the problems we are seeing are optical (i.e. imprinted on the PSL light) or electronic. My findings:

  1. 60 Hz line noise in PMC REFL and PMC TRANS is heavily dependent on whether I connect cables between the measuring PDs and Acromag ADC or not - but even with the Acromag cable disconnected, the 60 Hz RIN is HUGE - 10 mVpp out of 670 mV DC, and lines are much dirtier if you have connections to the SLOW ADCs. Measurement was made by looking at the time-domain signals on a battery powered Tektronix oscilloscope. See Attachment #1. I believe this line noise is higher it was. Cause is unknown to me at this point.
  2. The NPRO noise eater seems to function as advertised. The measured RIN with the noise eater enabled (our nominal operating condition) is in line with what the manual tells us it should be. See Attachment #2.
  3. There isn't strong evidence of excess frequency noise (measured with PLL) out to 100 kHz. I didn't measure the high-frequency part yet, but maybe I'm doing something wrong with the PLL setup which should be first corrected. See Attachments #3, #4.
  4. The beat note frequency between the free-running PSL and EX NPRO's is definitely slewing more than the quadrature sum of the advertised 1 MHz/min slewing per the manual.

Evidence:

Attachment #1: Time domain look at PMC Refl and Trans signals under various operating conditions. During this work, I took the chance to remove ~4 BNC T connectors that were connected on the PMC TRANS photodiode (Thorlabs). Now, there is one cable going to the Acromag ADC, and one going to the Oscilloscope used to monitor these signals. Any further T-ing can be done at the oscilloscope.

Attachment #2: RIN measurement of the NPRO light. I opted to place a Thorlabs PDA55 in the IR ALS pickoff light path. This is before the light sees the PMC. A DC block was inserted between the PDA55 and the AG4395 used to make the measurement. DC level of the PD output was 3.1 V into high-Z and I used half this value to normalize the measurement made by the 50-ohm input AG4395 into RIN units. The measurement was made with the PZT and slow temperature controls to the NPRO connected/disconnected, but I saw no significant difference. 

Attachment #3: Frequency noise measurement via PLL. This shows the loop transfer funtion for the PLL. Some details of the setup:

  • The beat note for locking the PLL was made between the PSL NPRO and the EX NPRO (output of the IR ALS BeatMouth). ~4dBm beatnote.
  • Local oscillator was sourced by a Marconi, f_carrier=33 MHz, RF level = +10dBm.
  • Level 7 Mixer and LB1005 controller from the mode-spectroscopy PLL setup.
  • PLL control signal routed to EX NPRO PZT via Heliax cable running along south arm. 
  • Why EX and not PSL or Marconi FM? Latter has limited range, ~1/10th of that offered by NPRO PZT. PSL PZT has a 2.9 Hz corner freq Pomona box. I could disconnect this for the purpose of PLL locking, but I thought it may be interesting to see if there’s any hints of the problem being electrical, by looking at PLL spectra with / without Pomona box. The expected delay due to cabling is only 400 ns, so not really a limiting factor for the PLL bandwidth.
  • LB 1005 settings:
    • PI corner = 3 kHz.
    • G = 2.30 (I could not increase this further - with the PSL+Lightwave NPRO PLL, we could achieve a UGF of ~60 kHz, but in this setup, I can't do much better than ~7kHz before the loop starts oscillating, not sure if the fact that the PZT actuation coefficient for the Innolight is ~5x lower than for the Lightwave is enough to explain this?).
    • LFGL = 90 dB.
  • Mixer output had a maximum value of 800 mVpp => PLL discriminant is 400 mV/rad.
  • The "eye fit" is just the transfer function of two poles at DC (one for frequency to phase conversion in the PLL and one for the LB1005 integrator), and a zero at 3kHz (PI corner). I scaled the gain till the "fit" and measurement lined up, and then used this model to undo the loop suppression of the error signal to extract the frequency noise without worrying about the frequency vector of the measurement being limited.
  • Once again, slow temperature control and PZT controls to the PSL NPRO were disconnected so this measurement was made with two free-running NPROs.

Attachment #4: Frequency noise measurement via PLL. This shows the frequency noise. I've overlaid the expected frequency noise between 2 free-running NPROs, model used is in the text box in the plot. There isn't strong evidence of excess high frequency noise in this measurement. The fact that the "LB 1005 input terminated" trace is below all the others supports the hypothesis that I'm measuring real frequency noise. The bump around a few kHz could indicate some gain peaking?

However, I'm unable to find good agreement between the measured frequency noise using the error point and the control point. For the former, I used the PLL discriminant mentioned above of 400 mV/rad, and undid the loop suppression, and for the latter I used a PZT discriminant of 1.7 MHz/V. However, there is still a constant scale difference between these two traces. So I'm doing something wrong?

Next steps:

  1. More interpretation of the PLL measurement results required.
  2. Measure the PLL error signal spectrum to higher frequencies using the AG4395. 
  3. ???

I've not disturbed the PLL setup in case anyone else wants to repeat these measurements, but I have restored the normal electrical connections to the PSL PZT and temperature control.

Some other activity:

  1. Alignment into the PMC was tweaked.
  2. NPRO laser pump current was increased from 1.9 A to 2.0 A.
  3. PMC servo gain was changed from +18 to +17 to prevent the servo from oscillating.
Attachment 1: consolidatedOscopeScreenCaps.pdf  246 kB  Uploaded Sun Jun 23 09:55:38 2019  | Hide | Hide all
consolidatedOscopeScreenCaps.pdf
Attachment 2: RINcomp.pdf  129 kB  Uploaded Sun Jun 23 09:55:51 2019  | Hide | Hide all
RINcomp.pdf
Attachment 3: PLL_OLTF.pdf  42 kB  Uploaded Sun Jun 23 09:56:35 2019  | Hide | Hide all
PLL_OLTF.pdf
Attachment 4: PLLnoise.pdf  40 kB  Uploaded Sun Jun 23 10:13:46 2019  | Hide | Hide all
PLLnoise.pdf
ELOG V3.1.3-